OpenStudy (anonymous):

What is the standard form equation of the circle given by the information below? Center: (0,0) Radius: Square root of 21

OpenStudy (anonymous):

A. \[x ^{2}+y ^{2}=21\]

OpenStudy (anonymous):

B\[x ^{2}+(y+21)^{2}\]

OpenStudy (anonymous):

c.\[(x+21)^{2}+y ^{2}=21\]

OpenStudy (anonymous):

D.\[x ^{2}+y ^{2}=\sqrt{21}\]

OpenStudy (anonymous):

the standard form is (x^2-h^2)+(y^2-k^2)=r^2

OpenStudy (anonymous):

thats not one of the options @joshhfers

OpenStudy (anonymous):

So when you plug them in they are (x^2-0)+(y^2-0)=square root of 21

OpenStudy (anonymous):

which is the same as x^2+y^2=and in the radius the square roots cancels.leaving it as 21